OpenStudy (anonymous):

A complete sentence explaining which method of solving would be best used based on the value of the discriminant. First Discriminant = 12.2i

OpenStudy (anonymous):

@SolomonZelman @Compassionate @juanpabloJR @nincompoop @ganeshie8 @texaschic101

OpenStudy (anonymous):

If you could help me through this that would be great!! Thank you

OpenStudy (anonymous):

I don't know what to do?

OpenStudy (anonymous):

@precal @paki @mathmale @Compassionate @Nurali

OpenStudy (anonymous):

12.2i sorry

OpenStudy (mathmale):

A complete sentence explaining which method of solving would be best used based on the value of the discriminant. First Discriminant = 12.2i Aiden: Hello! Note that the discriminant is never 12.2 i. No discriminant contains 'i'. Try again, please? Clarify what you want to know.

OpenStudy (anonymous):

\[\sqrt{-9.55^2-4(-1.91)(-7.64)}\] \[\sqrt{-91.2025-58.3696}\] \[\sqrt{-149.5721}\] That's my math....

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

OK. Notice you have no 'i' here yet. Your radical can be re-written as\[\sqrt{(-1)(149.9721)}\] which, in turn, can be re-written as \[\sqrt{-1}*\sqrt{\sqrt{?}} \sqrt{149.9721}=i \sqrt{149.9721}.\] Ayden, are you able to identify the "discriminant" in this example?

OpenStudy (mathmale):

Hint: Look under the radical sign in \[\sqrt{-149.5721}\] What you see UNDER the radical is your "discriminant."

OpenStudy (anonymous):

so my discriminant is -149.5721

OpenStudy (mathmale):

Your job, Ayden, is to explain how to write the solutions to the quadratic equation whose discriminant is -149.5721. The question you've posted is not worded very clearly (which is, of course, not your fault).

OpenStudy (mathmale):

Yes, Ayden, your discriminant is -149.5721. OpenStudy was down, which is why this response on my part is coming so late.

OpenStudy (mathmale):

I have to get off my computer, but will be on OpenStudy again later today. Best wishes to you.